OpenStudy (anonymous):

The area of the base of a prism is 55 mm2. The perimeter of the base is 30 mm. The height of the prism is 7 mm. What is the surface area of the prism? A. 445 mm2 B. 320 mm2 C. 265 mm2 D. 92 mm2

OpenStudy (anonymous):

 This is a prism right?

OpenStudy (anonymous):

it does not show

OpenStudy (anonymous):

well I got an answer and I can show you my method.

OpenStudy (anonymous):

x+y = 30, xy =55 x(30-x)=55 \[30x-x ^{2}=55\] \[x ^{?}-30x+55=0\] \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\rightarrow \frac{ 30\pm \sqrt{(-30)^{2}-4(1)(55)} }{ 2(1) }\]

OpenStudy (anonymous):

so it B

OpenStudy (anonymous):

x=28.04 or 1.96

OpenStudy (anonymous):

oh kk thanx

OpenStudy (anonymous):

you got the answer? I was still trying to see which one

OpenStudy (anonymous):

I guess you already know because you have x and y from the quadratic formula since they are both positive values. use the smaller length for the base for the triangle

OpenStudy (anonymous):

umm yeah it is B 320 mm2

OpenStudy (anonymous):

hmm I didn't get that. Not sure how you did.

OpenStudy (anonymous):

surface area should be a sum of all the faces: